OpenStudy (anonymous):

A mans salary goes from 16,000 to 19,000 What percentage raise has he gotten?

OpenStudy (anonymous):

Percent increase = (increase / original amount) x 100 What is the increase?

OpenStudy (anonymous):

i don't get it. at all .

OpenStudy (anonymous):

The increase in your case is just the increase in his salary. It goes up from 16000 to 19000. That's an increase of ______

OpenStudy (anonymous):

it goes up by 4 thousand...?

OpenStudy (anonymous):

3 my bad.

OpenStudy (anonymous):

correct. the original amount or old amount is where he started: 16000

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

so plug all of this into the formula i gave you in the first post

OpenStudy (anonymous):

i do not understand your formula !

OpenStudy (anonymous):

\[\frac{ increase }{ original amount }\times 100\]

OpenStudy (anonymous):

You know the increase: 3000 You know the original amount: 16000

OpenStudy (anonymous):

i still don't get it. i'll just get it wrong .

OpenStudy (anonymous):

\[\frac{ 3000 }{ 16000 }\times100\]

OpenStudy (anonymous):

Do you know how to put that in your calculater? 3000 divided by 16000 times 100 =

OpenStudy (mathstudent55):

\(\bf {percent ~change = \dfrac{\color{blue}{new~amount} - \color{red}{old~amount}}{\color{red}{old~amount}} \times 100 }\) \( \bf{ \color{blue}{new ~amount = 19,000 }} \) \( \bf{\color{red}{old ~amount = 16,000}} \) \(\bf {percent ~change = \dfrac{\color{blue}{19000} - \color{red}{16000}}{\color{red}{16000}} \times 100 }\) \(\bf {percent ~change = \dfrac{3000} {\color{red}{16000}} \times 100 }\) \(\bf {percent ~change = 18.75\% } \)